area a is 20m bigger than area b. 3/4 of area b is equal to half of area a. how big is area b? and mrs. collo is 1 1/2 taller than mrs. mole. 2/3 of mrs. mole's height is equal to the height of mrs. collo. how tall is mrs. mole?
1
Expert's answer
2010-10-13T11:21:34-0400
Let's construct and solve a system of equations:
1.
a = b + 20
3/4 b = 1/2a

b= 2/3 a;
a = 2/3 a + 20;
1/3 a = 20;
a = 60.
b = 2/3 * 60 = 40.

2. 1 1/2 = 3/2
M = C + 3/2;
2/3 M = C.

M =2/3 M +3/2;
1/3 M = 3/2;
M = 9/2 = 4.5;
C = M - 3/2 = 9/2 - 3/2 = 6/2 = 3.

Answers: b = 40; the height of mrs. Mole = 4.5